### v1

# ProLLM: Protein Chain-of-Thoughts Enhanced LLM for Protein-Protein Interaction Prediction

Abstract
The prediction of protein-protein interactions (PPIs) is crucial for understanding biological functions and diseases. Previous machine learning approaches to PPI prediction mainly focus on direct physical interactions, ignoring the broader context of nonphysical connections through intermediate proteins, thus limiting their effectiveness. The emergence of Large Language Models (LLMs) provides a new opportunity for addressing this complex biological challenge. By transforming structured data into natural language prompts, we can map the relationships between proteins into texts. This approach allows LLMs to identify indirect connections between proteins, tracing the path from upstream to downstream. Therefore, we propose a novel framework ProLLM that employs an LLM tailored for PPI for the first time. Specifically, we propose Protein Chain of Thought (ProCoT), which replicates the biological mechanism of signaling pathways as natural language prompts. ProCoT considers a signaling pathway as a protein reasoning process, which starts from upstream proteins and passes through several intermediate proteins to transmit biological signals to downstream proteins. Thus, we can use ProCoT to predict the interaction between upstream proteins and downstream proteins. The training of ProLLM employs the ProCoT format, which enhances the model’s understanding of complex biological problems. In addition to ProCoT, this paper also contributes to the exploration of embedding replacement of protein sites in natural language prompts, and instruction fine-tuning in protein knowledge datasets. We demonstrate the efficacy of ProLLM through rigorous validation against benchmark datasets, showing significant improvement over existing methods in terms of prediction accuracy and generalizability. Our results highlight the potential of LLMs to transform the field of PPI, serving as a robust potential tool for various categories of biological and medical research. The code is available at: https://github.com/MingyuJ666/ProLLM.
1 Introduction
Protein-protein interactions (PPIs) play an essential role in various biological processes of all living organisms, which are crucial for biomedical, genetic, and pharmaceutical research. Thus, numerous experimental methods have been proposed for PPI detection, such as yeast two-hybrid [1] and quantitative proteomics methods [2]. However, wet-lab methods for PPI prediction are often time-consuming and labor-intensive, highlighting the need for more precise and efficient computational tools.
In recent years, computational biology has developed rapidly. Methods such as the Convolutional Neural Network (CNN) and Graph Neural Network (GNN) have become powerful tools for studying protein interaction. CNN-based approaches like TAG-PPI [3] typically use pre-trained embedding models to convert the amino acid sequences in proteins into continuous vector representations, and then employ one-dimensional convolutional neural networks to extract features from the vectors for subsequent PPI tasks. Although CNN methods have shown some effectiveness in PPI prediction, they still have limitations due to their fixed receptive fields and the lack of well-defined spatial relationships in protein sequences, which limit the accuracy and interpretability of the predictions. GNN-based methods such as GNN-PPI [4] treat proteins as nodes and their relationships as edges, constructing a network composed of proteins, which better captures protein relationships and interactions, and outperforms CNNs in predicting protein interactions. However, while GNNs can effectively utilize network structural information, they are less effective at learning protein chains than transformer-based models  [5]. Furthermore, GNNs cannot fully capture and understand the relationships and dynamic changes in the signal passing process in real organisms, restricting their development for protein prediction  [6].
Figure 1: Illustration of the ProLLM Framework. We fine-tuning ProLLM under Human, SHS27K, SHS148K, and STRING datasets, enabling it to solve various PPI related tasks with the structure information purely described by natural language.
Following the GNN and CNN methods, Large Language Models have also been applied to this PPI area, such as ProBert [7] and ProteinLM [8]. As long as these models can obtain a protein representation, we can use the direct cosine similarity of the representation or train an MLP to perform PPI prediction. However, these methods still cannot capture the chain relationships between proteins, such as the signaling pathways. Besides, previous literature only used LLMs as a feature extractor. Recently, using LLM as a link predictor has shown that it can better capture relational information between nodes in knowledge graph tasks and its performance surpasses traditional GNN baselines [9, 10, 11]. Therefore, it is promising to introduce LLM for protein-protein interaction (PPI) tasks, since the most important biological signal for PPI tasks is the chain relationships of proteins, i.e., the signaling pathways.
To bridge the gap, we propose ProLLM, with its key ideas illustrated in Figure 1, and the difference between the existing method and our ProLLM shown in Figure 2. Existing methods only focus on the single protein-protein interaction, overlooking the application of protein links to predict PPI in signaling pathways. Instead, we employ a large language model to learn the law of signal pathways and adapt the LLM to directly predict the type of interaction between proteins.
The signaling pathway addresses the traditional method’s ignorance of global, non-physical connections between proteins. Signaling pathways typically start with an upstream protein that sends a biological signal through several intermediates to a downstream protein, hence requiring consideration of the cumulative effect of multiple protein interactions. This series of interactions form sequential chains. Therefore, we propose Protein Chain of Thought (ProCoT) to overcome the limitation in understanding signaling pathways and protein functions. ProCoT is a data format that simulates the signal transduction process using a thought-chain approach, thereby enabling the prediction of protein interactions in signaling pathway problems. CoT can express the thinking process step by step to form a reasoning chain [12], while ProCoT extends this principle further into the protein-related domain to simulate protein signaling pathways, giving LLMs a deeper insight into protein.
Figure 2: The difference between the existing method and our method in PPI prediction. Existing method focus on the property of upstream protein and downstream protein, our method focus on signaling pathway-like connection.
Additionally, our approach addresses the issue of poor protein comprehension in LLMs by replacing the standard language model embedding with embedding infused with protein information. When we process the protein name in the prompt, we replace its original embedding by ProtTrans [13], because its embedding contains protein’s structual information. We also perform the instruction fine-tuning on the protein knowledge dataset to infuse protein domain knowledge into LLMs. Following these steps, the LLM acquires a robust ability to reason about the direct relationships between proteins, as demonstrated in Figure 1. It can provide answers to questions about protein relationships, which play a significant role in biological research.
Our contributions are summarized as follows:
  • To the best of our knowledge, we are the first to explore the PPI prediction problem as a natural language processing problem. The proposed Protein Chain of Thought (ProCoT) is a novel method for understanding complex multi-step protein interactions, i.e., those present in signaling pathways.
  • We propose embedding replacement and instruction fine-tuning on our model to enhance its ability to understand and reason about proteins. This also provides our model with rich background knowledge of protein sequences and protein interactions before training.
  • Experiments on four widely used PPI datasets (i.e., Human, SHS27K, SHS148K, and STRING) demonstrate that ProLLM outperforms graph-based methods such as GNN-PPI and SemiGNN-PPI. It also has better performance than LLM-based methods like InstructGLM. The micro-F1 scores on these 4 datasets are 91.05, 85.32, 87.66, 89.21, respectively.
2 Related Work
Protein-protein Iteractions
Protein-protein interactions (PPIs) are indispensable for all living organisms, and so many efforts have been made for PPI detection up to now. Yeast two-hybrid (Y2H) assays [14], synthetic lethality [15], quantitative proteomics [16], and mass spectrometry [17], are widely used for identifying PPIs. To be specific, Y2H assays explore binary PPIs in living cells, offering insight into protein functions and regulations, although labor-intensive and limited in genomic coverage. Synthetic lethality identifies essential gene pair interactions by revealing compensatory relationships when both are inactivated. Meanwhile, quantitative proteomics, especially through mass spectrometry, illuminates the dynamic nature of the interactome under various conditions. Although informative, these methods require significant labor, time, and resources.
Traditional Machine Learning Models for PPI Prediction
In the realm of machine learning, sequence-based approaches include Shen’s SVM method [18], which uses a 3-mer count vector from protein sequences as features and groups 20 amino acids into seven classes to handle synonymous mutations and reduce feature space dimensionality. SVM-based methods [19] and the ensemble model PCA-EELM (Principal Component Analysis-Ensemble Extreme Learning Machine) [20] utilize various types of protein sequence information for PPI prediction. In the domain of deep learning, DeepPPI [21] extracts a multitude of features from protein sequences and employs a dual deep neural network structure for feature fusion and prediction. Sun et al. [22] introduced a PPI predictor based on a stacked autoencoder, emphasizing the importance of sample balance. DPPI [23] and TAGPPI [3] further extend the application of convolutional neural networks and integrate text convolution networks with graph representation learning to enhance the accuracy of PPI predictions. GNNs have significantly advanced PPI predictions, improving our understanding of biological mechanisms. GNN-PPI [4] enhances inter-novel-protein prediction accuracy by utilizing protein relationships and a new evaluation approach. PT-GNN [24] integrates diverse data for link prediction, learning node features from sequence and structure. DeepRank-GNN [25] offers a modular, Python-packaged framework for GNN-based interaction pattern predictions. HIGH-PPI [26] introduces a hierarchical graph learning model for effective PPI prediction and molecular detail extrapolation. Geometric GNNs excel in modeling spatial intricacies, enhancing biomolecule prediction accuracy. Geo-PPI [27] utilizes self-supervised learning for geometric protein structure representations, excelling in detailing protein interactions. mmCSM-PPI [28] captures multifaceted features for mutation impact predictions on protein interactions. MAPE-PPI [29] defines the microenvironment of amino acid residues in proteins and encodes it into discrete codes, which can capture the connections between different microenvironments, enhancing the prediction accuracy of PPI.
Large Language Model for PPI prediction
Recent advances in large language models, such as BERT [30], GPT [31], LLaMA [32], and T5 [33], have significantly advanced the field of Natural Language Processing (NLP) to new heights. These models, having been trained on extensive textual corpora, exhibit exceptional capabilities across a diverse range of NLP applications. Inspired by LLMs, Protein Large Language Models(PLMs) pre-trained on large-scale protein sequences have emerged, such as ESM [34], ProtTrans [13] and ProteinBert [13]. PLMs provide a better representation of protein sequences by converting the protein sequences into the form of high-dimensional numerical vectors, known as embedding. With the protein sequences captured by the PLMs, the performances on diverse downstream tasks, such as structure prediction [35], subcellular localization prediction [36], single peptide prediction [37] and N-linked glycosylation sites prediction [38], have been transformed. It can be expected that PLMs will assist in PPI prediction tasks. ProtLLM [10] utilizes a dynamic protein mounting mechanism, a protein-as-word language modeling approach, and the InterPT dataset for pre-training, enabling it to handle complex inputs and achieve superior performance on various proteins-related tasks. However, ProtLLM is used for general protein tasks, and it is not used for PPI tasks exactly. The methodology of training these LLMs to convert text inputs to desired text outputs positions them as particularly advantageous for tasks such as generative link prediction [9, 11]. In such tasks, the model is tasked with inferring and generating the relationship between two entities based on provided textual cues. Moreover, the extensive pre-training phase enables LLMs to exhibit a remarkable capacity for generalization. This capacity allows them to effectively tackle and respond to tasks or prompts that were not explicitly covered during their initial training [39]. In addition to the outlined capabilities, the inherent flexibility and generalization potential of LLMs suggest that their applicability extends well beyond the conventional boundaries of NLP tasks [40]. Specifically, their proficiency in generalizing from expansive pre-training sessions paves the way for their application in fields like bioinformatics and complex network analysis.
3 Proposed ProLLM Framework
In this section, we will introduce the implementation details of ProLLM, a framework designed to transform protein interaction data into ProCoT format natural language descriptions in order to simulate protein signaling pathways. By translating the structure and relationships of proteins into natural language, we effectively transform protein interaction problems into natural language processing (NLP) tasks. To enhance the understanding of proteins by ProLLM, we directly integrate the protein vectors generated from ProtTrans [13] into our model to replace the original word embedding in the protain name’s place. This approach allows our model to understand and utilize the biological attributes of proteins more accurately when predicting protein interactions. The ProLLM process is shown in Figure 3.
Figure 3: The process of ProLLM. Sector 1: Transfer the original protein data into ProCoT format of natural language that indicates the signaling pathways between proteins; Sector 2: Replace protein information embeddings with natural language embeddings to enhance the model’s understanding of proteins; Sector 3: Inject knowledge about protein function; Sector 4: Fine-tuning on the ProCoT format dataset in Sector 1.
3.1 ProCoT
3.1.1 Protein Data to Natural Language in CoT Format
Transformation of the original protein data into a ProCoT natural language is a critical step, this kind of natural language type that represents the relationships of proteins is shown as Definition 1.
Definition 1
(Protein Interaction) Given the set $\mathcal{P}$ representing all proteins under consideration in the study, we define the interaction between any two proteins $p_{n},p_{m}\in\mathcal{P}$ with the specific type interaction $R_{i}$ interaction set $\mathcal{R}$. This relationship is encapsulated as a tuple $(p_{n},p_{m},R_{i})$.
Definition 2
(Protein Signaling Pathway) In cellular signal transduction, receptor protein $\mathcal{R}\text{c}$ is the beginning of the signal pathway, and it can interact with downstream signaling proteins $\mathcal{S}\text{p}$. $\mathcal{S}\text{p}$ is responsible for continuing to transmit signals to effector protein $\mathcal{E}\text{p}$ inside the cell. After the $\mathcal{E}\text{p}$ response the signal, it will trigger specific biological effects and activate secondary messengers $\mathcal{M}$. $\mathcal{M}$ may be the end of this signal level, or it may open second signal hierarchy. If in the second signal level, the $\mathcal{M}$ will replace the $\mathcal{R}\text{c}$ in the initial signal hierarchy as the starting point. The signal will continue to propagate from $\mathcal{R}\text{c}$ through new signaling proteins, effector proteins, and then generate new secondary messengers. After this stage, the propagation of the signal repeats the structure of the second signal level until the signal is interrupted.
We aim to enhance the model’s understanding of biological signaling pathways, enhancing it’s ability to learn and reason interactions within complex signal networks. We create prompts in the ProCoT (Protein Chain of Thought) format to incrementally decompose the signal transduction process like Definition 2, simulating real pathways of signal propagation. Specifically, we clarify the structure knowledge within the signaling pathway, such as signaling proteins and effector proteins. We then design rules for signal transmission across different levels to simulate the iterative process of signal transduction in proteins. Finally, we also design signal interruptions to simulate the continuity of protein transmission in real life.
We use hard code to convert protein interaction information as Definition 1 in dataset into ProCoT format natural language to imitate protein signaling pathways as Definition 2. Figure 6 in the appendix is an example of ProCoT. The answer to this is "The relationship is activation".
3.1.2 Training on ProCot Format Dataset
In biology, co-expression and co-localization refer to the phenomenon in which proteins that are often expressed or located together in the cell tend to participate in the same or interconnected biological processes [41]. Thus, biologists frequently use known protein information to infer unknown protein interactions. Based on the principles of co-expression and co-localization in biology, we have formulated our training strategy. After constructing our ProCoT training data using DFS, we will obtain a circular protein interaction as in Figure 4. We mask the relationship between the initial and final proteins in the signaling pathway and let our model predict this relationship.
Figure 4: The fine-tuning process of ProCoT. Within the first dashed box, solid lines between proteins represent the signaling pathway, and the dashed lines connecting the head and tail proteins indicate the masked interaction. Our model will predict the type of masked interaction.
3.1.3 Why ProCot Works: Biological Intuition behind ProCot:
In this section, we will explore ProCoT through the lenses of biology and AI. (1) Simulating Biological Signaling Pathways: The biological signaling pathway is a series of ordered, rule-based interaction processes. During the process, the output of each step serves as the input for the next step, forming a highly organized information transmission chain. This pattern is consistent with the methods of handling serialized information in natural language processing (NLP). Additionally, Flan-T5’s self-attention blocks can capture the indirect relationships between proteins that are distant and do not have direct interactions. It is crucial for understanding biological signal transduction processes that involve multiple steps and complex intermediate links. (2) Signal Interruption Mechanism: This mechanism mimics protein adaptability. Because the signaling between proteins needs to be constantly interrupted to ensure the cell’s accurate response to external changes. This mechanism we designed can satisfy the complex feedback mechanisms and regulatory networks within the cell. Overall, our ProCoT design follows biological principles.
3.2 Enhancing Protein Sequence and Function Comprehension
In our methodology, we enhance the model’s understanding of protein sequence by replacing the T5 embedding with the ProTrans embedding vectors, and also perform instruction fine-tuning to enable the model to learn protein functions.
3.2.1 Embedding Replacement
We use a new embedding mechanism facilitated by the ProTrans [7] model to perform embedding replacement. ProTrans is a large-scale protein model capable of transforming protein sequences into embedding vectors with the biophysical and structural features of proteins.
Definition 3
(Embedding Replacement) Given the protein id $P_{\text{id}}$, we can query its protein sequence $P_{\text{seq}}$ by the professional bioinformatics tool Ensembl BioMart. $P_{\text{seq}}$ will be the input of ProTrans and ProTrans will output a $1\times 1024$-dimensional vector $P_{\text{emb}}$.
We add all protein IDs in the protein dataset and their corresponding embeddings to the T5 vocabulary. Because ProtTrans is trained based on T5-large, the process of adding to the vocabulary does not involve any change in dimensions, and there is no information loss.
The replaced embedding vectors from ProtTrans can add a lot of prior knowledge about the intrinsic patterns and biophysical features of proteins to our model, better applying it to subsequent protein interaction prediction tasks.
3.2.2 Instruction Finetuning
Mol-Instructions (Mol) [42] dataset is applied to conduct instruction fine-tuning on our model. Mol is a comprehensive instruction dataset designed for the field of biomolecules that aims to address the limitations of LLMs in biomolecular research. We use Mol to teach our model with knowledge of protein functions.
Definition 4
(Description of Mol) The content of the Mol dataset can be defined as follows: instructional text $I$ related to LLMs queries, an input amino acid sequence $S$ that includes essential protein information, and metadata $M$ that sheds light on vital details like the protein’s subcellular location, its primary function, and its participation in biological processes, followed by a corresponding output $O$ that serves as the response to $I$.
To facilitate instruction-based fine-tuning, we convert these objects into prompt-answer pairs $(P,A)$. This is shown as Figure 7 in the appendix. We convert the entire Mol dataset into a prompt-answer format and using these prompt-answer pairs for instruction fine-tuning of our model(ProLLM). This part can enhance the model’s predictive performance on protein-related tasks.
4 Experiments
In this section, we present the experimental results to answer the following research questions (RQs).
  • RQ1 - How does the performance of the proposed ProLLM framework compare to other baselines in terms of PPI prediction accuracy and generalizability?
  • RQ2 - How do different LLM backbones (e.g., Flan-T5-base and LLaMA-7b) affect the performance of the proposed ProLLM framework?
  • RQ3 - What are the contributions of each component to the PPI prediction performance of the ProLLM framework?
4.1 Experimental Settings
Datasets:
We undertake comprehensive evaluations on a trio of public Protein-Protein Interaction (PPI) datasets: Human [3], STRING [43], SHS27k, and SHS148k [44]. In these four datasets, we employ DFS dataset partitioning techniques. By prioritizing depth, DFS can effectively capture the step-by-step signal transmission and the hierarchical signal level of protein signaling pathways. Each dataset is split into training, validation, and testing sets, maintaining a proportion of 70%, 10%, and 20% respectively.
Baselines:
Earlier studies on predicting PPI are not pre-trained, they do not have prior knowledge about the proteins. Hence, we choose SVM [45], DPPI [23], DNN-PPI [46], PIPR [44], GNN-PPI [4] and SemiGNN-PPI [47] as the baseline without pretraining. Furthermore, we explore how protein pre-training can influence the PPI, we include ProBERT [7], SM-1b [48], GearNet-Edge [49], and KeAP [50] as pre-trained baselines. MAPE-PPI [29] and InstrucGLM [9] have two versions: one with pre-training and one without pre-training.
Evaluation Metrics:
We selected the micro-F1 score [51] as our evaluation metric because the PPI dataset exhibits class imbalance, making the F1 score a very relevant reference. Note that micro-F1 is widely used in the protein-protein interaction task [52]. Each dataset follows a 70% training, 10% validation, and 20% testing data split. Subsequently, we will choose different random seeds for training and testing, conducting a total of 10 tests. The mean micro-F1 score across these trials will serve as the definitive measure of model performance, with the accompanying standard deviation reflecting variability in different experimental runs.
4.2 Comparative Experiment (RQ1)
We compare the performance of ProLLM with other baselines (w/ and w/o additional pre-training data) on four datasets in Table 1. Note that here we use Flan-T5-large [33] as the backbone model. Based on the results, we can make three important observations: (1) Our method outperforms other baselines without pre-training. Although InstructGLM is also LLM based, it lags behind ProLLM; (2) As for the models pre-trained on protein dataset, they cannot achieve the performance of ProLLM without prior knowledge; (3) ProLLM outperforms GearNet-Edge, KeAP and MAPE-PPI, although they utilize a significantly larger dataset comprised of structural and knowledge graph data for pre-training than the Mol dataset.
Method Pre-training Dataset Human SHS27k SHS148k STRING
SVM - 61.28$\pm$1.28 53.07$\pm$5.16 58.59$\pm$0.07 64.59$\pm$0.03
DPPI - 54.19$\pm$0.78 46.12$\pm$3.02 52.03$\pm$1.18 66.82$\pm$0.29
DNN-PPI - 61.72$\pm$1.30 54.34$\pm$1.30 58.42$\pm$2.05 64.94$\pm$0.93
PIPR - 62.72$\pm$0.50 57.80$\pm$3.24 63.98$\pm$0.76 67.45$\pm$0.30
GNN-PPI - 78.61$\pm$1.38 66.52$\pm$5.26 75.34$\pm$1.54 84.28$\pm$0.89
SemiGNN-PPI - 80.79$\pm$1.40 69.25$\pm$3.91 77.62$\pm$1.08 84.85$\pm$0.65
MAPE-PPI - 82.13$\pm$1.47 72.04$\pm$3.46 80.45$\pm$1.12 86.48$\pm$0.52
InstructGLM - 81.35$\pm$2.04 70.01$\pm$3.75 75.35$\pm$1.98 84.15$\pm$1.85
ProLLM(Flan-T5-large) - 87.32$\pm$1.93 83.87$\pm$3.54 85.13$\pm$1.86 87.12$\pm$1.68
ProBERT BFD 79.58$\pm$0.76 68.85$\pm$3.18 74.76$\pm$1.21 83.82$\pm$0.49
EMS-1b UniRef50 81.48$\pm$1.02 70.69$\pm$3.40 79.64$\pm$1.93 85.21$\pm$0.76
KeAP ProteinKG25 82.31$\pm$0.71 72.38$\pm$2.96 80.20$\pm$1.26 86.58$\pm$0.41
GearNet-Edge AlphaFoldDB 82.87$\pm$1.02 72.06$\pm$3.56 79.84$\pm$1.65 85.96$\pm$1.01
MAPE-PPI CATH4.2 83.64$\pm$1.22 73.21$\pm$2.97 81.78$\pm$1.24 87.23$\pm$0.35
InstructGLM Mol dataset 85.71$\pm$2.01 81.63$\pm$3.44 83.41$\pm$1.78 85.25$\pm$1.72
ProLLM(Flan-T5-large) Mol dataset 91.05$\pm$1.63 85.32$\pm$3.29 87.66$\pm$1.68 89.21$\pm$1.45
Table 1: The micro-F1 of different methods (w/o and w/ additional pre-training data) on different datasets, where bold and underline denote the best and second best metrics, respectively. Higher micro-F1 denotes better performance.
4.3 Influence of Different Backbones (RQ2)
We choose Flan-T5-base, Flan-T5-large, Flan-T5-XL [33] and LLaMA-7b [32] models as the backbone for ProLLM. We report the micro-F1 performance comparison as in Table 2. We should replace the embedding in Flan-T5 by ProtTrans and ProtTrans is a pretrained model based on Flan-T5-Large. Therefore, the embedding generated by ProtTrans will match Flan-T5-large better. Additionally, despite having more model parameters, LLaMA-v1-7b exhibits worse lower micro-F1 scores in four datasets compared to lighter models: Flan-T5-base, Flan-T5-large, and Flan-T5-XL. Furthermore, the greater standard deviation of LLaMA-v1-7b highlights its instability.
Method Pre-training Dataset Human SHS27k SHS148k STRING
ProLLM-Flan-T5-base - 82.62$\pm$2.01 76.37$\pm$4.11 78.19$\pm$1.82 80.63$\pm$1.97
ProLLM-Flan-T5-large - 87.32$\pm$1.93 83.87$\pm$3.54 85.13$\pm$1.86 87.12$\pm$1.68
ProLLM-Flan-T5-XL - 84.32$\pm$2.65 79.51$\pm$3.69 81.73$\pm$2.21 82.94$\pm$1.93
ProLLM-LLaMA-v1-7b - 81.75$\pm$4.46 74.29$\pm$6.26 77.08$\pm$3.59 79.41$\pm$2.87
ProLLM-Flan-T5-base Mol dataset 87.92$\pm$2.07 82.84$\pm$3.97 85.07$\pm$1.87 87.11$\pm$1.88
ProLLM-Flan-T5-large Mol dataset 91.05$\pm$1.63 85.32$\pm$3.29 87.66$\pm$1.93 89.21$\pm$1.45
ProLLM-Flan-T5-XL Mol dataset 89.16$\pm$2.41 83.47$\pm$3.12 86.13$\pm$1.97 87.97$\pm$1.78
ProLLM-LLaMA-v1-7b Mol dataset 87.08$\pm$4.72 82.39$\pm$6.31 84.61$\pm$3.53 86.71$\pm$2.61
Table 2: The micro-F1 score of ProLLM on different backbones. Higher micro-F1 denotes better performance. Bold and underline denote the best and second-best metrics, respectively.
4.4 Ablation Study (RQ3)
In our ablation study (Table 3), we evaluated the impact of different configurations on PPI prediction. The configurations are as follows: ProLLM w/o ProCoT: This setup shuffles ProCoT data, disrupting signaling pathways to mimic the model’s performance without pathway understanding. It limits the model’s ability to learn from signaling sequences. ProLLM relies on memorizing fixed relations instead of reasoning through intermediate protein relationships; ProLLM w/o Embedding Replacement: It evaluates the effect of protein-specific embedding features; ProLLM w/o Instruction Fine-tuning: This setup examines the model’s capability to predict protein-protein interactions without the application of instruction fine-tuning on Mol dataset; ProLLM w/o Embedding and Instruction Fine-tuning: This configuration tests the model’s performance without utilizing both protein-specific embedding features and instruction fine-tuning on Mol dataset.
In our ablation study, we have identified that the presence of Procot has the most significant impact on the performance of ProLLM. Our experiments revealed that introducing Procot led to substantial improvements in the performance of the model. Additionally, we explored other techniques such as embedding replacement and instruction fine-tuning on the Mol dataset. While these approaches did show some positive effects on the model’s performance, their impact was found to be comparatively smaller when compared to the influence of ProCoT.
ProCoT Embedding Replacement Instruction Fine-tuning Dataset
Human SHS27k SHS148k STRING
83.87$\pm$1.25 76.35$\pm$1.47 79.12$\pm$2.64 82.68$\pm$1.36
87.32$\pm$1.93 83.87$\pm$3.54 85.13$\pm$1.86 87.12$\pm$1.68
88.53$\pm$1.45 83.17$\pm$2.25 85.94$\pm$3.39 87.62$\pm$1.68
78.32$\pm$2.65 72.42$\pm$1.95 74.10$\pm$1.27 77.85$\pm$1.54
91.03$\pm$1.63 85.36$\pm$3.29 87.64$\pm$1.93 89.28$\pm$1.45
Table 3: Ablation study. The metric here is micro-F1. Where bold denote the best metrics. Higher micro-F1 denotes better performance.
5 Conclusions and Future Work
In this paper, we propose ProLLM, a novel framework that leverages LLMs for protein-protein interaction prediction by representing protein data in natural language formats. Our key contributions include: 1) ProCoT (Protein Chain of Thought) to convert multi-step protein signaling pathways to natural language prompts, and the design of ProCoT can reflect the actual protein signaling passing within a biological organism. Additionally, the format of ProCoT is sequential, which is a type of information that LLMs are good at processing. 2) Integration of protein-specific embeddings from ProtTrans, and 3) Instruction fine-tuning on protein knowledge datasets. Through extensive experiments on four PPI datasets, ProLLM significantly outperformed existing graph-based and language model methods in prediction accuracy and generalizability. By unifying language models with structured biological data, our work opens up new possibilities for driving discoveries in computational biology, drug discovery, and broader scientific domains.